What is the first step to solve w:
4/5=w/20

Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

What is the first step to solve w:
4/5=w/20
------
Multiply both sides by 20 to get:
w = 20(4/5)
----
w = 4*4 = 16
======================
Cheers,
Stan H.

Answer by richwmiller(17219)

The only step necessary is multiply both sides by 20
4/5*20=w/20*20
16=w